Consider a simple bond has a coupon payment at the end of 1 year. The Face value of the bond is $1,000 and the stated interest rate on the bond is R=0.05 (5%) per year. If you buy this bond today, and then the interest rate on a newly issued bonds of this type goes up to R= 0.10 (10%) right after you buy yours, the most you could sell your bond for right now would be a. $916.67 b. $900 c. 954.45 d. $909.09 E. $950
Answer #1

So, here the face value of the bond is “$1,000” and the coupon payment is “5%”, => 1,000*0.05 = $50. Now, if interest rate increases to “10”, => the value of the bond is given by.

=> B = 50/1.1 + 1000/1.1 = 45.4545 + 909.0909 = 954.54, => here the correct answer is “C”,